OpenStudy (anonymous):

Suppose a curve C such that its slope is dy/dx= -x/y. If C passes through the point (-4, 3), answer the following. a) Find an equation for the curve. b) Find the points where the curve has a vertical tangent line.

OpenStudy (anonymous):

okay! i solved part a. the equations is sqrt(25-x) = y(x) how do i do part b?

ganeshie8 (ganeshie8):

you should get \(y(x) = \pm \sqrt{25-x^2}\)

ganeshie8 (ganeshie8):

its a circle of radius 5

ganeshie8 (ganeshie8):

use below for part b : vertical tangents occur when slope is undefined

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

would it be 25?

ganeshie8 (ganeshie8):

what would be 25 ?

OpenStudy (anonymous):

x=25 is where a tangent line is?

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

\(\large \frac{dy}{dx} = \frac{-x}{y}\) slope has "undefined" value when the denominator, \(y = 0\)

ganeshie8 (ganeshie8):

so the x values where the tangent is vertical can be found by setting y = 0 : \(\sqrt{25-x^2} = 0\) solve \(x\)

OpenStudy (anonymous):

oh! okay thanks!

ganeshie8 (ganeshie8):

np :)